OpenStudy (anonymous):

I have to verify that this is true using trig identities sec^4(x)−sec^2(x)=tan^4(x)+tan^2(x)

OpenStudy (anonymous):

consider left hand side sec^4(x)−sec^2(x) = {sec^2(x)}^2-(1+tan^2(x)) = {1+tan^2(x)}^2-1-tan^2(x) =1+tan^4(x)+2tan^2(x)-1-tan^2(x) =tan^4(x)+tan^2(x)=right hand side hence proved!!!!!!

OpenStudy (anonymous):

Thank you!!! Wow .. that was too simple.

OpenStudy (anonymous):

:)